Jefferson Lab HALL A Logbook for April 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

  • 40327 04/30/00 23:49 Hyde-Wright shift summary
  • 40318 04/30/00 18:07 Hyde-Wright bpm
  • 40315 04/30/00 17:42 Hyde-Wright spot
  • 40314 04/30/00 17:40 Adaq fpp stripchart
  • 40305 04/30/00 12:38 M Jones access to look at FPP eff problem
  • 40301 04/30/00 11:45 E. Christy IOC fan dead (replaced)
  • 40297 04/30/00 10:46 mitchell statistics so far Q^2=2.6
  • 40290 04/30/00 08:12 hubt shift summary
  • 40286 04/30/00 06:29 R. Michaels Raster Plots
  • 40285 04/30/00 06:14 R. Michaels BPM snapshots
  • 40282 04/30/00 05:03 R. Michaels Online ESPACE (using ET)
  • 40275 04/30/00 01:48 hubt FPP eff
  • 40274 04/30/00 01:05 R. Michaels spot is far off, per instructions ?
  • 40271 04/30/00 00:38 hubt FPP gas strip
  • 40270 04/30/00 00:36 hubt magnet chart
  • Saturday

  • 40263 04/29/00 18:44 S. Dieterich Target IOC reboot
  • 40260 04/29/00 17:39 S. Dieterich BCM coefficients
  • 40258 04/29/00 17:10 S. Dieterich Target IOC reboot
  • 40257 04/29/00 16:55 A. Gasparian controled access
  • 40253 04/29/00 15:24 Ulmer Statistics projection Q^2=2.6
  • 40252 04/29/00 13:58 tforest Harp scan
  • 40241 04/29/00 11:43 tforest BCM calibration
  • 40238 04/29/00 09:33 tforest beam delivery
  • 40199 04/29/00 01:11 R. Michaels status, test runs
  • Friday

  • 40179 04/28/00 14:11 Olivier end of access
  • 40178 04/28/00 13:23 Olivier Controlled access for Compton
  • 40177 04/28/00 09:56 Olivier FPP wires fixed
  • 40172 04/28/00 06:08 R. Michaels lots of test runs
  • 40141 04/28/00 04:33 R. Michaels status and wierd computer problem.
  • 40091 04/28/00 02:00 Tony Trigger Setup and Checked
  • 40065 04/28/00 01:34 Tony FPP wires dead
  • Thursday

  • 40047 04/27/00 23:14 jpchen cosmic run 1855 started
  • 40046 04/27/00 23:11 jpchen Gas Alarm fixed, All HVs on.
  • 40043 04/27/00 22:18 jpchen Gas alarm
  • 40042 04/27/00 22:16 jpchen Chnage from control access to beam permit
  • 40041 04/27/00 22:10 jpchen spectrometer settings
  • 40040 04/27/00 21:28 jpchen Spectrometers are set for Q2=2.6 GeV2
  • Wednesday

  • 40025 04/26/00 15:43 R. Michaels "ER disconnect" problem fixed
  • Tuesday

  • 40011 04/25/00 15:04 R. Michaels New disk situation
  • 40002 04/25/00 14:48 R. Michaels new DAQ disks, tests
  • Monday

  • 39993 04/24/00 14:18 R. Michaels adaqs2 shutdown for awhile
  • 39992 04/24/00 08:43 mitchell statistics so far
  • 39991 04/24/00 08:21 S. Strauch magnet screen at beginning of maintenance
  • 39989 04/24/00 08:20 S. Strauch Owl Shift Summary
  • 39988 04/24/00 07:35 S. Strauch FPP gas - strip chart
  • 39976 04/24/00 00:00 Z. Chai Swing Shift Summary
  • Sunday

  • 39972 04/23/00 22:24 Eugene Moller results
  • 39958 04/23/00 16:12 M JOnes Shift summary
  • 39951 04/23/00 14:21 M Jones bcm situation
  • 39925 04/23/00 08:05 S. Strauch Shift summary (Owl shift)
  • 39921 04/23/00 07:01 S. Strauch Strip tool of FPP gas
  • 39910 04/23/00 00:00 Kumbartzki swing shift summary
  • Saturday

  • 39897 04/22/00 15:56 M JOnes SHift Summary
  • 39896 04/22/00 15:49 M Jones fpp gas log
  • 39893 04/22/00 14:40 M JOnes Access to hall
  • 39892 04/22/00 13:36 M JOnes rebooted target ioc
  • 39890 04/22/00 12:10 M Jones rebooted hac ioc
  • 39885 04/22/00 10:53 mitchell statistics so far He Q^2=1.6
  • 39882 04/22/00 09:01 M JOnes BCM still down.
  • 39879 04/22/00 08:31 M JOnes IOC16 message
  • 39877 04/22/00 08:02 tony OWL Shift Summary
  • 39870 04/22/00 04:27 tony BCMs down
  • 39867 04/22/00 03:34 tony screenshot of fpp gas
  • 39866 04/22/00 03:23 tony post lightning strike
  • 39865 04/22/00 03:13 tony magnet screen grab
  • Friday

  • 39855 04/21/00 23:46 Kumbartzki swing shift summary
  • 39839 04/21/00 16:05 gilman day shift summary
  • 39825 04/21/00 09:24 mitchell statistics so far
  • 39824 04/21/00 09:23 R. Michaels re: CODA error msg (not)
  • 39820 04/21/00 08:01 olivier Owl shift summary
  • 39819 04/21/00 07:57 Olivier CODA error message
  • 39812 04/21/00 06:25 Olivier Reset Q1 display
  • 39808 04/21/00 04:42 tony FPP stripchart
  • 39803 04/21/00 02:12 Olivier Reset Q1 display
  • 39802 04/21/00 01:22 tony Straw Efficiency (rear chmb.)
  • 39798 04/21/00 00:00 Kumbartzki swing shift summary
  • Thursday

  • 39784 04/20/00 16:13 gilman day shift summary
  • 39781 04/20/00 15:54 Dieterich Target CT 96
  • 39780 04/20/00 15:13 Dieterich Target IOC died
  • 39772 04/20/00 10:43 Saha Harp scans 5 and 6
  • 39768 04/20/00 10:10 mitchell statistics so far: He Q^2 = 1.6
  • 39762 04/20/00 08:00 Z. Chai Owl Shift Summary
  • 39751 04/20/00 03:28 Z. Chai FPP Gas Strip Plot
  • 39742 04/20/00 00:02 R Roche Shift Summary
  • Wednesday

  • 39736 04/19/00 21:49 R Roche FPP gas strip chart
  • 39732 04/19/00 18:55 Eugene Moller target problem
  • 39730 04/19/00 16:20 kuss harp scan
  • 39728 04/19/00 15:53 kuss harp scan
  • 39727 04/19/00 14:16 Eugene Moller results
  • 39726 04/19/00 12:10 dieterich FPP straw efficiency
  • 39725 04/19/00 12:07 Dieterich FPP straw eff.
  • 39724 04/19/00 11:36 kuss Moller starts
  • 39716 04/19/00 08:09 Z. Chai Owl Shift Summary
  • 39712 04/19/00 06:06 Z. Chai L1_T1_Aft_heater_R
  • 39711 04/19/00 05:57 S. Strauch FPP Gas Strip Chart
  • 39694 04/19/00 00:07 Finn Shift summary
  • Tuesday

  • 39692 04/18/00 23:56 R Roche Loop 1 Aft Temp
  • 39681 04/18/00 17:57 M Jones Time estimate for Q2=1.0 4He
  • 39679 04/18/00 17:49 R Roche Loop 1 Aft Temp
  • 39670 04/18/00 16:01 Saha Shift Summary, 0800-1600, April 18, 2000
  • 39667 04/18/00 15:24 jpchen Q1 tripped
  • 39653 04/18/00 09:51 jpchen beam spot size
  • 39645 04/18/00 08:07 M. Rvachev Shift summary
  • 39644 04/18/00 08:05 M. Rvachev fpp gas log
  • Monday

  • 39633 04/17/00 22:09 gayou magnet screen
  • 39632 04/17/00 22:08 gayou fpp gas screen
  • 39631 04/17/00 22:03 gayou Camera does not move
  • 39629 04/17/00 22:02 finn hadron arm spectrometer moved to 45.743 deg
  • 39630 04/17/00 22:00 gayou Magnet screen
  • 39628 04/17/00 20:55 R. Michaels halog old archives
  • 39624 04/17/00 11:10 Ulmer Q^2=1.6 1H statistics
  • 39608 04/17/00 08:04 M. Rvachev Shift summary
  • 39596 04/17/00 06:25 tony FPP gas stripchart
  • 39560 04/17/00 00:04 finn swing shift summary 4/16/2000
  • Sunday

  • 39540 04/16/00 19:56 Ulmer Beam position and Harps
  • 39531 04/16/00 18:36 Finn Q1 trip Run 1646
  • 39530 04/16/00 18:28 mitchell improved estimate of statistics: using deadtime
  • 39521 04/16/00 16:13 Ulmer Q^2=1.6 1H statistics
  • 39517 04/16/00 15:46 Saha Harp scanners 5 and 6 OK
  • 39516 04/16/00 15:43 Ulmer Compton data to exclude
  • 39514 04/16/00 15:29 Saha Arc Energy Measurment, 5 pass, 4.235 GeV
  • 39513 04/16/00 14:01 Eugene Moller results
  • 39512 04/16/00 11:45 mitchell database serverup again
  • 39502 04/16/00 09:28 ulmer FPP gas strip chart
  • 39495 04/16/00 08:03 M. Rvachev Shift summary
  • 39492 04/16/00 07:58 M. Rvachev fpp gas log
  • 39446 04/16/00 00:08 M Jones fpp gas log
  • Saturday

  • 39443 04/15/00 23:54 glashausser shift summary
  • 39429 04/15/00 21:41 R. Michaels Disk logjam solved
  • 39420 04/15/00 20:29 M Jones copying files to the MSS
  • 39419 04/15/00 20:26 R Roche moved files
  • 39418 04/15/00 20:23 glashausser mag screen for 1.6 GeV^2
  • 39404 04/15/00 17:59 M Jones Yesterday's access fixed FPP problem
  • 39403 04/15/00 17:50 R. Michaels adaqs3, adaqfs disks
  • 39402 04/15/00 17:47 R. Michaels stage disk 100% full again
  • 39397 04/15/00 16:09 ransome change to carbon
  • 39394 04/15/00 15:36 ransome disk space out again
  • 39388 04/15/00 15:05 ransome data files copied to gain space
  • 39353 04/15/00 11:57 ransome ER1 disconnect
  • 39342 04/15/00 10:47 Adaq magnet screen
  • 39321 04/15/00 08:04 M. Rvachev Shift summary
  • 39317 04/15/00 07:43 R. Michaels Disk logjam
  • 39229 04/15/00 00:05 perdrisat 16-24 shift continued
  • Friday

  • 39222 04/14/00 23:34 glashausser magnet screen
  • 39209 04/14/00 22:38 perdrisat log for 16-24 shift
  • 39199 04/14/00 21:43 baker hydrogen running start
  • 39173 04/14/00 18:31 mitchell Target change
  • 39172 04/14/00 18:28 mitchell ep hardware problems
  • 39171 04/14/00 18:17 baker harp scan problem persists
  • 39169 04/14/00 14:50 ransome shift summary
  • 39168 04/14/00 12:49 sarty Total T5 trigs for He4, Q^2=1.0
  • 39163 04/14/00 09:41 ransome runs off center
  • 39161 04/14/00 09:24 ransome beam off for retune
  • 39156 04/14/00 08:20 Gayou Owl shift summary
  • 39139 04/14/00 05:06 gayou FPP gas strip charts
  • 39140 04/14/00 05:06 boswell magnet screen
  • Thursday

  • 39115 04/13/00 23:57 ransome shift summary
  • 39106 04/13/00 22:05 R. Michaels spot, run 1456
  • 39105 04/13/00 21:57 Keppel total events
  • 39098 04/13/00 20:33 ransome raster problem?
  • 39077 04/13/00 16:19 sarty NEW how to use the MYSQL database to get trigger total
  • 39076 04/13/00 16:12 glashausser shift summary
  • 39069 04/13/00 15:21 sarty how to use the MYSQL database to get trigger total
  • 39066 04/13/00 14:47 M Jones problem with ch1 fixed
  • 39055 04/13/00 12:25 sarty sum T5 trigs-to-date at Q^2=1.0
  • 39048 04/13/00 11:06 glashaus magnet screen
  • 39041 04/13/00 10:22 M Jones Access to hadron hut
  • 39036 04/13/00 08:20 S. Strauch Owl shift summary
  • 39006 04/13/00 00:09 baker short shift summary (SSS)
  • Wednesday

  • 38995 04/12/00 22:06 Eugene Moller results
  • 38992 04/12/00 21:39 sarty spot size & centering
  • 38990 04/12/00 20:59 Jack Ethane mass flow controller
  • 38989 04/12/00 18:27 M Jones E_miss resolution
  • 38988 04/12/00 17:28 M Jones estimate of running time at Q2=1 4He
  • 38986 04/12/00 15:14 M Jones Slight problem with fpp
  • 38978 04/12/00 08:19 S. Strauch Shift summary
  • 38974 04/12/00 07:16 Dieterich Target Loop 2
  • 38961 04/12/00 04:23 S. Strauch Magnet screen for 4He(e,e'p) @ 1.0 GeV^2
  • 38956 04/12/00 03:24 S. Strauch FFP HV trip
  • Tuesday

  • 38931 04/11/00 19:47 sarty ep ion chamber masked off
  • 38929 04/11/00 15:59 mitchell shift summary
  • 38906 04/11/00 10:00 R. Michaels no more 'accepted events' entries
  • 38901 04/11/00 07:30 Meekins New trigger count
  • 38883 04/11/00 05:34 Meekins Bad Runs
  • 38882 04/11/00 05:32 Meekins FPP HV trip
  • 38823 04/11/00 00:15 Nilanga Emiss resolution
  • Monday

  • 38816 04/10/00 23:44 fissum fpp gas
  • 38815 04/10/00 23:43 fissum status sheets, shift end
  • 38814 04/10/00 23:39 fissum run 1329 hung
  • 38806 04/10/00 23:15 fissum reset HRSh Q1 display
  • 38802 04/10/00 22:40 fissum run 1324 hung
  • 38797 04/10/00 22:29 fissum reset HRSh Q1 display
  • 38786 04/10/00 21:20 fissum run 1316 hung
  • 38775 04/10/00 20:20 fissum fpp gas
  • 38773 04/10/00 20:18 fissum status sheets, shift middle
  • 38772 04/10/00 20:16 fissum FPP chamber currents OK
  • 38767 04/10/00 19:42 fissum run 1309 hung
  • 38762 04/10/00 19:28 fissum /data4, run 1307
  • 38753 04/10/00 18:40 fissum run 1303 hung
  • 38748 04/10/00 18:18 fissum /data4, run 1301
  • 38741 04/10/00 17:59 fissum no Moeller
  • 38736 04/10/00 17:33 fissum new sample spectra
  • 38733 04/10/00 17:18 fissum /data4
  • 38726 04/10/00 16:25 fissum fpp gas
  • 38725 04/10/00 16:24 fissum status sheets shift beginning
  • 38722 04/10/00 16:12 fissum 4/10 Swing shift goals
  • 38709 04/10/00 15:27 Ulmer Q^2=0.5 4He stats as of 15:20
  • 38706 04/10/00 15:05 adev Accepted_Events_Run_1285
  • 38704 04/10/00 14:55 adev Accepted_Events_Run_1284
  • 38700 04/10/00 14:43 adev Accepted_Events_Run_1283
  • 38697 04/10/00 14:32 adev Accepted_Events_Run_1282
  • 38694 04/10/00 14:17 adev Accepted_Events_Run_1281
  • 38691 04/10/00 14:13 adev Accepted_Events_Run_1280
  • 38689 04/10/00 14:02 adev Accepted_Events_Run_1279
  • 38686 04/10/00 13:51 adev Accepted_Events_Run_1278
  • 38682 04/10/00 13:39 adev Accepted_Events_Run_1277
  • 38679 04/10/00 13:34 Ulmer FPP gas flow strip chart
  • 38678 04/10/00 13:29 adev Accepted_Events_Run_1276
  • 38675 04/10/00 13:14 adev Accepted_Events_Run_1275
  • 38671 04/10/00 12:58 adev Accepted_Events_Run_1273
  • 38669 04/10/00 12:52 adev Accepted_Events_Run_1272
  • 38668 04/10/00 12:49 R. Michaels double pulsing of E-arm trigger
  • 38665 04/10/00 12:43 R. Michaels Trigger change, coinc time
  • 38664 04/10/00 12:40 adev Accepted_Events_Run_1271
  • 38659 04/10/00 12:22 adev Accepted_Events_Run_1270
  • 38658 04/10/00 12:11 adev Accepted_Events_Run_1269
  • 38654 04/10/00 11:56 adev Accepted_Events_Run_1268
  • 38651 04/10/00 11:25 Nilanga emiss resolution
  • 38650 04/10/00 11:22 Ulmer E-arm trigger double pulsing
  • 38649 04/10/00 11:17 Ulmer Coincidence gate reduced
  • 38646 04/10/00 10:25 Ulmer Window guards, con'td.
  • 38645 04/10/00 10:10 Ulmer Scatt. chmbr. shield in the way!
  • 38644 04/10/00 09:44 jpchen One Clarification
  • 38643 04/10/00 09:42 adev Accepted_Events_Run_1266
  • 38640 04/10/00 09:13 adev Accepted_Events_Run_1265
  • 38638 04/10/00 09:00 adev Accepted_Events_Run_1264
  • 38635 04/10/00 08:48 adev Accepted_Events_Run_1259
  • 38634 04/10/00 08:47 adev Accepted_Events_Run_1259
  • 38633 04/10/00 08:47 adev Accepted_Events_Run_1259
  • 38632 04/10/00 08:46 adev Accepted_Events_Run_1263
  • 38629 04/10/00 08:30 adev Accepted_Events_Run_1262
  • 38626 04/10/00 08:16 adev Accepted_Events_Run_1261
  • 38617 04/10/00 07:43 adev Accepted_Events_Run_1258
  • 38614 04/10/00 07:23 adev Accepted_Events_Run_1257
  • 38611 04/10/00 07:12 adev Accepted_Events_Run_1256
  • 38608 04/10/00 07:02 adev Accepted_Events_Run_1255
  • 38605 04/10/00 06:49 adev Accepted_Events_Run_1254
  • 38602 04/10/00 06:39 adev Accepted_Events_Run_1253
  • 38599 04/10/00 06:25 adev Accepted_Events_Run_1252
  • 38597 04/10/00 06:23 adev Accepted_Events_Run_1251
  • 38593 04/10/00 06:12 adev Accepted_Events_Run_1250
  • 38590 04/10/00 06:01 adev Accepted_Events_Run_1249
  • 38587 04/10/00 05:48 adev Accepted_Events_Run_1248
  • 38584 04/10/00 05:36 adev Accepted_Events_Run_1247
  • 38581 04/10/00 05:24 adev Accepted_Events_Run_1246
  • 38577 04/10/00 05:09 adev Accepted_Events_Run_1244
  • 38576 04/10/00 05:08 adev Accepted_Events_Run_1243
  • 38572 04/10/00 04:58 adev Accepted_Events_Run_1242
  • 38569 04/10/00 04:47 adev Accepted_Events_Run_1241
  • 38566 04/10/00 04:35 adev Accepted_Events_Run_1240
  • 38563 04/10/00 04:25 adev Accepted_Events_Run_1239
  • 38560 04/10/00 04:12 adev Accepted_Events_Run_1237
  • 38557 04/10/00 04:06 adev Accepted_Events_Run_1236
  • 38554 04/10/00 03:56 adev Accepted_Events_Run_1235
  • 38551 04/10/00 03:43 adev Accepted_Events_Run_1234
  • 38548 04/10/00 03:32 adev Accepted_Events_Run_1233
  • 38544 04/10/00 03:18 adev Accepted_Events_Run_1231
  • 38543 04/10/00 03:14 adev Accepted_Events_Run_1230
  • 38539 04/10/00 03:04 adev Accepted_Events_Run_1229
  • 38536 04/10/00 02:51 adev Accepted_Events_Run_1228
  • 38533 04/10/00 02:41 adev Accepted_Events_Run_1227
  • 38530 04/10/00 02:30 adev Accepted_Events_Run_1226
  • 38527 04/10/00 02:20 adev Accepted_Events_Run_1225
  • 38524 04/10/00 02:06 adev Accepted_Events_Run_1224
  • 38521 04/10/00 02:03 adev Accepted_Events_Run_1223
  • 38518 04/10/00 01:49 adev Accepted_Events_Run_1222
  • 38515 04/10/00 01:37 adev Accepted_Events_Run_1221
  • 38512 04/10/00 01:26 adev Accepted_Events_Run_1220
  • 38509 04/10/00 01:13 adev Accepted_Events_Run_1219
  • 38506 04/10/00 01:02 adev Accepted_Events_Run_1218
  • 38503 04/10/00 00:51 adev Accepted_Events_Run_1217
  • 38500 04/10/00 00:40 adev Accepted_Events_Run_1216
  • 38496 04/10/00 00:28 adev Accepted_Events_Run_1215
  • 38494 04/10/00 00:21 adev Accepted_Events_Run_1214
  • 38490 04/10/00 00:09 adev Accepted_Events_Run_1213
  • 38488 04/10/00 00:09 Higinbotham ep energy measurement
  • Sunday

  • 38485 04/09/00 23:56 adev Accepted_Events_Run_1212
  • 38483 04/09/00 23:50 Higinbotham Shift Summary
  • 38482 04/09/00 23:44 adev Accepted_Events_Run_1211
  • 38479 04/09/00 23:30 adev Accepted_Events_Run_1210
  • 38476 04/09/00 23:24 adev Accepted_Events_Run_1209
  • 38472 04/09/00 23:15 adev Accepted_Events_Run_1208
  • 38470 04/09/00 23:13 fissum beam conditions checked
  • 38467 04/09/00 22:58 fissum magnets are ready
  • 38466 04/09/00 22:42 fissum fpp information record
  • 38465 04/09/00 21:55 Jack Gasmaster alert for e- arm.
  • 38464 04/09/00 21:27 fissum stubborn HRSe dipole
  • 38463 04/09/00 20:23 folts cryo problems
  • 38462 04/09/00 18:29 Higinbotham Q1 and Target Problems
  • 38461 04/09/00 17:02 R. Michaels disk situation
  • 38460 04/09/00 17:01 Higinbotham Low Ethane
  • 38459 04/09/00 16:10 adev Accepted_Events_Run_1207
  • 38457 04/09/00 15:10 Ulmer H-arm Q1 smoking
  • 38456 04/09/00 14:07 M JOnes entry to hadron detector hut
  • 38455 04/09/00 13:23 ulmer magnet summary
  • 38454 04/09/00 13:22 ulmer energy strip charts - 2nd try
  • 38453 04/09/00 13:18 ulmer energy strip charts
  • 38452 04/09/00 13:17 kuss 15K target warm up
  • 38451 04/09/00 13:15 Ulmer 1H Q^2=0.5
  • 38450 04/09/00 13:02 adev Accepted_Events_Run_1206
  • 38446 04/09/00 12:42 adev Accepted_Events_Run_1204
  • 38443 04/09/00 12:34 adev Accepted_Events_Run_1203
  • 38440 04/09/00 12:24 adev Accepted_Events_Run_1202
  • 38437 04/09/00 12:19 adev Accepted_Events_Run_1201
  • 38434 04/09/00 12:12 adev Accepted_Events_Run_1200
  • 38431 04/09/00 12:03 adev Accepted_Events_Run_1199
  • 38428 04/09/00 11:55 adev Accepted_Events_Run_1198
  • 38425 04/09/00 11:48 adev Accepted_Events_Run_1197
  • 38424 04/09/00 11:48 adev Accepted_Events_Run_1192
  • 38423 04/09/00 11:48 adev Accepted_Events_Run_1192
  • 38421 04/09/00 11:43 adev Accepted_Events_Run_1195
  • 38419 04/09/00 11:40 adev Accepted_Events_Run_1196
  • 38414 04/09/00 11:07 R. Michaels mysterious absence of 1184
  • 38413 04/09/00 10:56 adev Accepted_Events_Run_1194
  • 38410 04/09/00 10:53 adev Accepted_Events_Run_1193
  • 38404 04/09/00 10:26 ulmer FPP gas flow
  • 38403 04/09/00 10:22 adev Accepted_Events_Run_1191
  • 38400 04/09/00 10:17 adev Accepted_Events_Run_1190
  • 38399 04/09/00 10:11 R. Michaels data disks full
  • 38396 04/09/00 09:42 adev Accepted_Events_Run_1189
  • 38394 04/09/00 09:40 adev Accepted_Events_Run_1185
  • 38391 04/09/00 09:28 adev Accepted_Events_Run_1188
  • 38387 04/09/00 09:14 adev Accepted_Events_Run_1186
  • 38379 04/09/00 08:43 adev Accepted_Events_Run_1184
  • 38378 04/09/00 08:41 adev Accepted_Events_Run_1183
  • 38375 04/09/00 08:35 adev Accepted_Events_Run_1182
  • 38372 04/09/00 08:26 adev Accepted_Events_Run_1181
  • 38367 04/09/00 08:14 ent shift summary
  • 38366 04/09/00 08:10 adev Accepted_Events_Run_1180
  • 38365 04/09/00 08:04 adev Accepted_Events_Run_1179
  • 38362 04/09/00 07:55 adev Accepted_Events_Run_1178
  • 38358 04/09/00 07:47 adev Accepted_Events_Run_1177
  • 38355 04/09/00 07:39 adev Accepted_Events_Run_1176
  • 38352 04/09/00 07:30 adev Accepted_Events_Run_1175
  • 38349 04/09/00 07:25 adev Accepted_Events_Run_1174
  • 38346 04/09/00 07:07 adev Accepted_Events_Run_1173
  • 38343 04/09/00 07:01 adev Accepted_Events_Run_1171
  • 38340 04/09/00 06:43 adev Accepted_Events_Run_1170
  • 38337 04/09/00 06:36 adev Accepted_Events_Run_1169
  • 38334 04/09/00 06:27 adev Accepted_Events_Run_1168
  • 38333 04/09/00 06:26 adev Accepted_Events_Run_1167
  • 38328 04/09/00 06:10 adev Accepted_Events_Run_1166
  • 38325 04/09/00 06:02 adev Accepted_Events_Run_1165
  • 38322 04/09/00 05:53 adev Accepted_Events_Run_1164
  • 38319 04/09/00 05:48 adev Accepted_Events_Run_1163
  • 38316 04/09/00 05:39 adev Accepted_Events_Run_1162
  • 38313 04/09/00 05:31 adev Accepted_Events_Run_1161
  • 38309 04/09/00 05:20 adev Accepted_Events_Run_1159
  • 38307 04/09/00 05:17 adev Accepted_Events_Run_1160
  • 38304 04/09/00 05:10 adev Accepted_Events_Run_1158
  • 38300 04/09/00 04:57 adev Accepted_Events_Run_1157
  • 38298 04/09/00 04:55 adev Accepted_Events_Run_1156
  • 38295 04/09/00 04:46 adev Accepted_Events_Run_1155
  • 38292 04/09/00 04:41 adev Accepted_Events_Run_1154
  • 38289 04/09/00 04:36 adev Accepted_Events_Run_1153
  • 38286 04/09/00 04:18 adev Accepted_Events_Run_1152
  • 38283 04/09/00 04:10 adev Accepted_Events_Run_1151
  • 38280 04/09/00 04:02 adev Accepted_Events_Run_1150
  • 38277 04/09/00 03:54 adev Accepted_Events_Run_1149
  • 38273 04/09/00 03:45 adev Accepted_Events_Run_1148
  • 38271 04/09/00 03:40 adev Accepted_Events_Run_1147
  • 38267 04/09/00 03:25 adev Accepted_Events_Run_1146
  • 38264 04/09/00 03:18 adev Accepted_Events_Run_1145
  • 38262 04/09/00 03:08 adev Accepted_Events_Run_1144
  • 38259 04/09/00 03:00 adev Accepted_Events_Run_1143
  • 38256 04/09/00 02:51 adev Accepted_Events_Run_1142
  • 38253 04/09/00 02:41 adev Accepted_Events_Run_1141
  • 38250 04/09/00 02:30 adev Accepted_Events_Run_1140
  • 38248 04/09/00 02:28 adev Accepted_Events_Run_1139
  • 38243 04/09/00 02:20 Meekins angle recording change for HRSe
  • 38244 04/09/00 02:19 adev Accepted_Events_Run_1138
  • 38240 04/09/00 02:12 adev Accepted_Events_Run_1137
  • 38236 04/09/00 02:03 adev Accepted_Events_Run_1136
  • 38233 04/09/00 01:47 adev Accepted_Events_Run_1134
  • 38232 04/09/00 01:45 adev Accepted_Events_Run_1133
  • 38229 04/09/00 01:37 adev Accepted_Events_Run_1132
  • 38225 04/09/00 01:29 adev Accepted_Events_Run_1131
  • 38222 04/09/00 01:19 adev Accepted_Events_Run_1128
  • 38219 04/09/00 01:14 adev Accepted_Events_Run_1130
  • 38217 04/09/00 01:12 adev Accepted_Events_Run_1129
  • 38212 04/09/00 01:00 adev Accepted_Events_Run_1127
  • 38210 04/09/00 00:56 adev Accepted_Events_Run_1126
  • 38206 04/09/00 00:41 adev Accepted_Events_Run_1125
  • 38203 04/09/00 00:29 adev Accepted_Events_Run_1123
  • 38200 04/09/00 00:23 adev Accepted_Events_Run_1122
  • 38198 04/09/00 00:12 adev Accepted_Events_Run_1121
  • 38195 04/09/00 00:00 adev Accepted_Events_Run_1120
  • 38193 04/09/00 00:00 Higinbotham Shift Summary and Accouting
  • Saturday

  • 38191 04/08/00 23:47 adev Accepted_Events_Run_1119
  • 38188 04/08/00 23:38 fissum sample spectra binder
  • 38187 04/08/00 23:33 adev Accepted_Events_Run_1118
  • 38184 04/08/00 23:22 adev Accepted_Events_Run_1117
  • 38181 04/08/00 23:09 adev Accepted_Events_Run_1116
  • 38178 04/08/00 22:57 adev Accepted_Events_Run_1115
  • 38175 04/08/00 22:45 adev Accepted_Events_Run_1114
  • 38172 04/08/00 22:32 adev Accepted_Events_Run_1113
  • 38169 04/08/00 22:15 adev Accepted_Events_Run_1112
  • 38166 04/08/00 22:01 adev Accepted_Events_Run_1111
  • 38163 04/08/00 21:47 adev Accepted_Events_Run_1110
  • 38160 04/08/00 21:41 Higinbotham Magnet Screen
  • 38159 04/08/00 21:35 adev Accepted_Events_Run_1109
  • 38154 04/08/00 21:17 adev Accepted_Events_Run_1108
  • 38152 04/08/00 20:27 Adaq triggers to date
  • 38151 04/08/00 20:04 adev Accepted_Events_Run_1107
  • 38149 04/08/00 19:53 adev Accepted_Events_Run_1106
  • 38147 04/08/00 19:47 Higinbotham Arc Energy (4 pass)
  • 38145 04/08/00 19:38 adev Accepted_Events_Run_1105
  • 38140 04/08/00 19:03 adev Accepted_Events_Run_1104
  • 38139 04/08/00 18:57 adev Accepted_Events_Run_1103
  • 38135 04/08/00 18:04 M Jones fpp gas chart
  • 38134 04/08/00 18:00 M Jones day shift summary
  • 38133 04/08/00 17:45 adev Accepted_Events_Run_1102
  • 38132 04/08/00 17:45 adev Accepted_Events_Run_1102
  • 38131 04/08/00 17:44 M Jones slight mistake in message 38130
  • 38130 04/08/00 17:37 M Jones estimate of run time for 1H
  • 38127 04/08/00 17:09 adev Accepted_Events_Run_1101
  • 38124 04/08/00 16:56 adev Accepted_Events_Run_1100
  • 38120 04/08/00 16:40 adev Accepted_Events_Run_1099
  • 38117 04/08/00 16:27 M Jones Changed hadron angle 63.12
  • 38116 04/08/00 16:27 adev Accepted_Events_Run_1098
  • 38113 04/08/00 16:17 X. jiang main control pannel at run 1098.
  • 38112 04/08/00 16:09 adev Accepted_Events_Run_1097
  • 38109 04/08/00 15:57 hubt coincidence event rate change
  • 38108 04/08/00 15:50 adev Accepted_Events_Run_1096
  • 38103 04/08/00 14:06 adev Accepted_Events_Run_1095
  • 38101 04/08/00 14:02 adev Accepted_Events_Run_1094
  • 38099 04/08/00 13:45 adev Accepted_Events_Run_1093
  • 38096 04/08/00 13:32 adev Accepted_Events_Run_1092
  • 38093 04/08/00 13:15 adev Accepted_Events_Run_1091
  • 38090 04/08/00 12:56 M JOnes beam back doing 1H
  • 38088 04/08/00 12:44 M Jones Fpp rear chamber eff
  • 38086 04/08/00 12:43 adev Accepted_Events_Run_1090
  • 38085 04/08/00 12:25 M Jones beam down
  • 38083 04/08/00 12:13 M Jones Front fpp chamber eff
  • 38082 04/08/00 11:30 adev Accepted_Events_Run_1089
  • 38080 04/08/00 10:57 M Jones fpp gas log
  • 38079 04/08/00 10:56 M JOnes carbon doors for run 1089
  • 38077 04/08/00 10:30 adev Accepted_Events_Run_1088
  • 38074 04/08/00 10:25 adev Accepted_Events_Run_1087
  • 38070 04/08/00 10:17 adev Accepted_Events_Run_1085
  • 38068 04/08/00 10:15 adev Accepted_Events_Run_1084
  • 38064 04/08/00 10:04 adev Accepted_Events_Run_1083
  • 38062 04/08/00 10:02 Ent shift summary
  • 38061 04/08/00 09:53 X. Jiang spot picture at run 1083.
  • 38060 04/08/00 09:48 adev Accepted_Events_Run_1082
  • 38056 04/08/00 09:25 adev Accepted_Events_Run_1081
  • 38054 04/08/00 09:19 adev Accepted_Events_Run_1080
  • 38050 04/08/00 08:52 adev Accepted_Events_Run_1079
  • 38048 04/08/00 08:39 X. jiang Magnet window at run 1079
  • 38046 04/08/00 08:25 adev Accepted_Events_Run_1078
  • 38042 04/08/00 07:57 adev Accepted_Events_Run_1077
  • 38040 04/08/00 07:27 ent beam spot/position
  • 38038 04/08/00 07:24 adev Accepted_Events_Run_1076
  • 38035 04/08/00 07:20 adev Accepted_Events_Run_1075
  • 38032 04/08/00 07:14 ent Wien filter saga
  • 38031 04/08/00 07:03 adev Accepted_Events_Run_1074
  • 38028 04/08/00 06:55 adev Accepted_Events_Run_1073
  • 38025 04/08/00 06:49 Eugene Moller results
  • 38024 04/08/00 01:47 Eugene Wien filter setting
  • 38023 04/08/00 01:45 Eugene Wrong polarity of Moller quads 1 and 3
  • Friday

  • 38021 04/07/00 23:58 Higinbotham Shift Summary
  • 38020 04/07/00 22:59 tony follow up to BPM mystery
  • 38019 04/07/00 22:28 adev Accepted_Events_Run_1072
  • 38016 04/07/00 22:26 fissum upstream BCM DVM swapped out
  • 38017 04/07/00 22:23 Higinbotham Strange Spot
  • 38014 04/07/00 21:58 adev Accepted_Events_Run_1071
  • 38011 04/07/00 21:52 fissum IOC 5 problems
  • 38010 04/07/00 21:48 adev Accepted_Events_Run_1070
  • 38007 04/07/00 21:07 fissum IOC 5 woes continue...
  • 38006 04/07/00 20:59 fissum IOC 5 error status
  • 38004 04/07/00 20:37 adev Accepted_Events_Run_1069
  • 38002 04/07/00 19:15 fissum IOC16 fault light
  • 38000 04/07/00 18:14 adev Accepted_Events_Run_1067
  • 37999 04/07/00 18:14 adev Accepted_Events_Run_1068
  • 37996 04/07/00 17:45 B. Hu Hadron Q1 display problem
  • 37995 04/07/00 17:15 M Jones Reboot ed BCM crate
  • 37993 04/07/00 16:27 M Jones Shift log
  • 37992 04/07/00 16:21 M JOnes log of fpp gas
  • 37991 04/07/00 14:22 M Jones remember after rebooting BCM crate
  • 37988 04/07/00 13:51 adev Accepted_Events_Run_1066
  • 37986 04/07/00 13:40 adev Accepted_Events_Run_1065
  • 37983 04/07/00 13:34 M Jones BCM crate controller rebooted
  • 37982 04/07/00 13:28 adev Accepted_Events_Run_1064
  • 37980 04/07/00 12:10 hubt Q1 tripped
  • 37978 04/07/00 10:45 adev Accepted_Events_Run_1063
  • 37975 04/07/00 10:32 M Jones Hadron Q1 tripped
  • 37974 04/07/00 09:47 adev Accepted_Events_Run_1062
  • 37972 04/07/00 09:26 x. Jiang H-arm Q2 reset.
  • 37971 04/07/00 09:01 M Jones electron dipole tripped
  • 37968 04/07/00 08:14 ent night shift
  • 37966 04/07/00 06:36 adev Accepted_Events_Run_1061
  • 37964 04/07/00 04:52 adev Accepted_Events_Run_1060
  • 37961 04/07/00 04:02 adev Accepted_Events_Run_1059
  • Thursday

  • 37958 04/06/00 22:45 Z. Chai Target Status
  • 37957 04/06/00 22:29 M Jones FPP status chambers 3 and 4
  • 37956 04/06/00 22:27 M Jones FPP status chambers 1 and 2
  • 37954 04/06/00 18:51 adev Accepted_Events_Run_1058
  • 37952 04/06/00 18:09 adev Accepted_Events_Run_1057
  • 37948 04/06/00 16:13 adev Accepted_Events_Run_1056
  • 37944 04/06/00 16:10 adev Accepted_Events_Run_1055
  • 37942 04/06/00 14:41 adev Accepted_Events_Run_1054
  • 37939 04/06/00 12:19 adev Accepted_Events_Run_1053
  • 37936 04/06/00 11:43 adev Accepted_Events_Run_1052
  • 37931 04/06/00 11:20 adev Accepted_Events_Run_1050
  • 37930 04/06/00 10:26 adev Accepted_Events_Run_1049
  • 37927 04/06/00 09:12 adev Accepted_Events_Run_1048
  • Wednesday

  • 37924 04/05/00 23:51 adev Accepted_Events_Run_1047
  • 37915 04/05/00 21:07 adev Accepted_Events_Run_1046
  • 37913 04/05/00 21:05 adev Accepted_Events_Run_1045
  • 37912 04/05/00 20:54 adev Accepted_Events_Run_1044
  • 37910 04/05/00 20:51 adev Accepted_Events_Run_1043
  • 37908 04/05/00 19:14 adev Accepted_Events_Run_1042
  • 37905 04/05/00 18:11 adev Accepted_Events_Run_1041
  • 37902 04/05/00 15:16 adev Accepted_Events_Run_1040
  • 37900 04/05/00 12:27 adev Accepted_Events_Run_1039
  • Tuesday

  • 37894 04/04/00 22:04 adev Accepted_Events_Run_1037
  • 37891 04/04/00 19:23 adev Accepted_Events_Run_1036
  • 37888 04/04/00 19:20 adev Accepted_Events_Run_1035
  • 37885 04/04/00 19:19 adev Accepted_Events_Run_1034
  • 37882 04/04/00 16:47 adev Accepted_Events_Run_1033
  • 37880 04/04/00 16:46 R. Michaels network problems
  • 37879 04/04/00 16:33 adev Accepted_Events_Run_1031
  • 37875 04/04/00 16:02 adev Accepted_Events_Run_1030
  • 37872 04/04/00 15:55 adev Accepted_Events_Run_1027
  • 37867 04/04/00 11:59 adev Accepted_Events_Run_1025
  • 37864 04/04/00 11:12 adev Accepted_Events_Run_1024
  • 37862 04/04/00 11:10 adev Accepted_Events_Run_1023
  • 37858 04/04/00 11:08 adev Accepted_Events_Run_1022
  • 37857 04/04/00 11:07 adev Accepted_Events_Run_1021
  • 37852 04/04/00 11:03 adev Accepted_Events_Run_1020
  • 37849 04/04/00 10:59 adev Accepted_Events_Run_1019
  • 37845 04/04/00 10:55 adev Accepted_Events_Run_1018
  • 37843 04/04/00 10:52 adev Accepted_Events_Run_1017
  • 37840 04/04/00 10:26 adev Accepted_Events_Run_1016
  • Monday

  • 37837 04/03/00 21:08 adev Accepted_Events_Run_1015
  • 37836 04/03/00 20:23 adev Accepted_Events_Run_1014
  • 37832 04/03/00 20:14 adev Accepted_Events_Run_1013
  • 37827 04/03/00 20:03 adev Accepted_Events_Run_1012
  • 37825 04/03/00 19:02 adev Accepted_Events_Run_1011
  • 37823 04/03/00 18:50 adev Accepted_Events_Run_1010
  • 37819 04/03/00 18:28 adev Accepted_Events_Run_1009
  • 37816 04/03/00 17:38 adev Accepted_Events_Run_1008
  • 37813 04/03/00 17:37 adev Accepted_Events_Run_1007
  • 37809 04/03/00 17:06 adev Accepted_Events_Run_1004
  • 37808 04/03/00 17:04 R. Michaels tests.........
  • 37806 04/03/00 17:00 adev Accepted_Events_Run_1003
  • 37802 04/03/00 16:58 adev Accepted_Events_Run_1002
  • 37801 04/03/00 16:57 R. M. test
  • 37800 04/03/00 08:08 Perdrisat shift summary
  • 37798 04/03/00 07:36 adev Accepted_Events_Run_3176
  • 37797 04/03/00 07:36 adev Accepted_Events_Run_3176
  • 37796 04/03/00 07:32 adev Accepted_Events_Run_3180
  • 37795 04/03/00 07:32 adev Accepted_Events_Run_3176
  • 37793 04/03/00 07:06 adev Accepted_Events_Run_3179
  • 37790 04/03/00 06:38 adev Accepted_Events_Run_3178
  • 37782 04/03/00 05:41 adev Accepted_Events_Run_3175
  • 37779 04/03/00 05:34 adev Accepted_Events_Run_3174
  • 37774 04/03/00 05:04 adev Accepted_Events_Run_3173
  • 37773 04/03/00 04:56 adev Accepted_Events_Run_3172
  • 37770 04/03/00 04:29 adev Accepted_Events_Run_3171
  • 37767 04/03/00 04:17 Perdrisat leadglass test shift summary
  • 37765 04/03/00 03:51 adev Accepted_Events_Run_3169
  • 37762 04/03/00 03:36 adev Accepted_Events_Run_3167
  • 37759 04/03/00 03:12 adev Accepted_Events_Run_3166
  • 37756 04/03/00 02:49 adev Accepted_Events_Run_3165
  • 37754 04/03/00 02:36 adev Accepted_Events_Run_3164
  • 37749 04/03/00 00:33 adev Accepted_Events_Run_3163
  • 37747 04/03/00 00:23 adev Accepted_Events_Run_3162
  • 37745 04/03/00 00:21 adev Accepted_Events_Run_3161
  • 37741 04/03/00 00:09 adev Accepted_Events_Run_3160
  • Sunday

  • 37737 04/02/00 23:59 adev Accepted_Events_Run_3159
  • 37734 04/02/00 23:31 adev Accepted_Events_Run_3158
  • 37731 04/02/00 22:57 adev Accepted_Events_Run_3157
  • 37729 04/02/00 22:41 adev Accepted_Events_Run_3156
  • 37726 04/02/00 22:29 adev Accepted_Events_Run_3155
  • 37722 04/02/00 20:55 adev Accepted_Events_Run_3154
  • 37718 04/02/00 20:33 adev Accepted_Events_Run_3153
  • 37716 04/02/00 20:27 adev Accepted_Events_Run_3152
  • 37712 04/02/00 20:18 adev Accepted_Events_Run_3151
  • 37709 04/02/00 20:04 adev Accepted_Events_Run_3150
  • 37707 04/02/00 19:58 adev Accepted_Events_Run_3149
  • 37705 04/02/00 19:34 R. Michaels timing adjustments (II)
  • 37703 04/02/00 18:43 adev Accepted_Events_Run_3148
  • 37701 04/02/00 18:40 R. Michaels trigger mods for Pedrisat test
  • 37700 04/02/00 18:34 adev Accepted_Events_Run_3147
  • 37697 04/02/00 18:28 adev Accepted_Events_Run_3146
  • 37693 04/02/00 16:25 mitchell ep energy measurement online result
  • 37692 04/02/00 16:20 Adaq e89044-denergy screen
  • 37691 04/02/00 16:19 M Jones e89044_beam screen and MEDM screen
  • 37689 04/02/00 16:15 M. Jones Shift Summary
  • 37688 04/02/00 14:55 Higinbotham Arc Energy (5 pass beam)
  • 37687 04/02/00 11:46 adev Accepted_Events_Run_3145
  • 37684 04/02/00 11:04 adev Accepted_Events_Run_3144
  • 37681 04/02/00 09:03 adev Accepted_Events_Run_3143
  • 37677 04/02/00 07:01 adev Accepted_Events_Run_3142
  • 37673 04/02/00 06:19 adev Accepted_Events_Run_3141
  • 37671 04/02/00 06:01 adev Accepted_Events_Run_3139
  • 37668 04/02/00 05:54 adev Accepted_Events_Run_3139
  • 37666 04/02/00 05:37 adev Accepted_Events_Run_3138
  • 37663 04/02/00 04:13 adev Accepted_Events_Run_3137
  • 37660 04/02/00 03:28 adev Accepted_Events_Run_3136
  • 37658 04/02/00 03:07 M. Rvachev Rebooted target IOC
  • 37656 04/02/00 03:02 adev Accepted_Events_Run_3135
  • 37653 04/02/00 00:26 gomez denergy
  • 37652 04/02/00 00:16 adev Accepted_Events_Run_3134
  • 37650 04/02/00 00:08 Eric Voutier Shift Summary
  • Saturday

  • 37647 04/01/00 23:47 adev Accepted_Events_Run_3133
  • 37645 04/01/00 23:44 adev Accepted_Events_Run_3132
  • 37641 04/01/00 23:22 Eric Voutier Magnet settings for Kin#301
  • 37640 04/01/00 23:11 adev Accepted_Events_Run_3131
  • 37638 04/01/00 23:02 adev Accepted_Events_Run_3130
  • 37632 04/01/00 05:40 adev Accepted_Events_Run_3128
  • 37630 04/01/00 05:32 R. Michaels tests finished. start cosmics run
  • 37627 04/01/00 05:30 adev Accepted_Events_Run_3127
  • 37626 04/01/00 05:28 R. Michaels trigger deadtime study
  • 37624 04/01/00 05:24 adev Accepted_Events_Run_3126
  • 37621 04/01/00 05:13 adev Accepted_Events_Run_3125
  • 37618 04/01/00 05:09 adev Accepted_Events_Run_3124
  • 37615 04/01/00 05:05 adev Accepted_Events_Run_3123
  • 37612 04/01/00 04:58 adev Accepted_Events_Run_3122
  • 37609 04/01/00 04:51 adev Accepted_Events_Run_3121
  • 37606 04/01/00 04:47 adev Accepted_Events_Run_3120
  • 37603 04/01/00 04:42 adev Accepted_Events_Run_3119
  • 37600 04/01/00 04:27 adev Accepted_Events_Run_3118
  • 37597 04/01/00 04:24 adev Accepted_Events_Run_3117
  • 37594 04/01/00 04:09 adev Accepted_Events_Run_3116
  • 37591 04/01/00 03:53 adev Accepted_Events_Run_3115
  • 37589 04/01/00 03:40 R. Michaels Sieve was in, now 6 msr.
  • 37586 04/01/00 03:27 adev Accepted_Events_Run_3114
  • 37585 04/01/00 03:27 R. Michaels background check
  • 37584 04/01/00 03:24 R. Michaels raster change
  • 37582 04/01/00 03:11 adev Accepted_Events_Run_3113
  • 37581 04/01/00 03:10 R. Michaels Elastic carbon runs 3112 and 3113
  • 37578 04/01/00 02:41 adev Accepted_Events_Run_3112
  • 37575 04/01/00 02:17 R. Michaels Cannot measure BPM pedestals
  • 37573 04/01/00 02:10 adev Accepted_Events_Run_3111
  • 37572 04/01/00 02:09 R. Michaels beam returned to (0,0)
  • 37569 04/01/00 02:06 adev Accepted_Events_Run_3110
  • 37568 04/01/00 02:04 R. Michaels (-1, 0) BPM
  • 37567 04/01/00 02:03 tony harp scan
  • 37564 04/01/00 02:00 tony Re-Entry for harp scan #768
  • 37563 04/01/00 02:00 adev Accepted_Events_Run_3109
  • 37562 04/01/00 01:59 R. Michaels bpm at (-1, +1)
  • 37560 04/01/00 01:50 tony scan
  • 37558 04/01/00 01:48 adev Accepted_Events_Run_3108
  • 37557 04/01/00 01:48 tony harp scan
  • 37556 04/01/00 01:47 R. Michaels beam (+1, 0)
  • 37555 04/01/00 01:46 tony harp scan
  • 37554 04/01/00 01:45 tony harp scan
  • 37552 04/01/00 01:43 tony harp scan
  • 37550 04/01/00 01:42 adev Accepted_Events_Run_3107
  • 37549 04/01/00 01:41 tony harp scan
  • 37548 04/01/00 01:41 R. Michaels bpm (+1, -1)
  • 37546 04/01/00 01:39 tony harp scan
  • 37544 04/01/00 01:37 adev Accepted_Events_Run_3106
  • 37543 04/01/00 01:37 R. Michaels bpm at (+1, +1)
  • 37542 04/01/00 01:35 tony harp scan
  • 37541 04/01/00 01:35 R. Michaels BeO spot comments
  • 37539 04/01/00 01:32 tony harp scan
  • 37538 04/01/00 01:29 adev Accepted_Events_Run_3105
  • 37536 04/01/00 01:28 R. Michaels beam (-1, -1)
  • 37535 04/01/00 01:26 tony harp scan
  • 37532 04/01/00 01:23 adev Accepted_Events_Run_3104
  • 37531 04/01/00 01:22 R. Michaels BPM at (-1, 0)
  • 37528 04/01/00 01:19 adev Accepted_Events_Run_3103
  • 37527 04/01/00 01:18 R. Michaels beam position (0, -1)
  • 37524 04/01/00 01:14 adev Accepted_Events_Run_3102
  • 37523 04/01/00 01:12 R. Michaels BPM calib: (0,+1)
  • 37520 04/01/00 01:07 adev Accepted_Events_Run_3101
  • 37519 04/01/00 01:07 R. Michaels HARP scan at (0,0)
  • 37517 04/01/00 00:50 R. Michaels owl run plan - no lead tests